OpenStudy (anonymous):

A particle is projected at an angle such that range and MAx height are same when again projected at same velocity but at different angle range is unchanged what is the ratio of the two time of flight?

OpenStudy (maheshmeghwal9):

Range & Maximum height are always equal only & only when body is thrown approximately at an angle of 76 degrees then i think ration is 1. but btw t is the answer?

OpenStudy (maheshmeghwal9):

t = wt*

OpenStudy (anonymous):

The options are (a) 2 (b) 3 (c) 2.5

OpenStudy (maheshmeghwal9):

wt is the different angle?

OpenStudy (anonymous):

since it is said that Range is same if one angle is x and other one is 90-x

OpenStudy (maheshmeghwal9):

but u haven't stated in the question that range is unchanged @Yahoo! :/

OpenStudy (maheshmeghwal9):

but where is this written that range is unchanged?????????????

OpenStudy (anonymous):

sorry for the mistake

OpenStudy (maheshmeghwal9):

np:)

OpenStudy (anonymous):

nw u read the question

OpenStudy (maheshmeghwal9):

ok! now it is much better:-)

OpenStudy (anonymous):

Range, Initial Velocity are same for both the projectiles but the angle is different? @Yahoo!

OpenStudy (anonymous):

Yup ....but that is second case

OpenStudy (anonymous):

There are two cases?! What's the first case?

OpenStudy (anonymous):

Range = Maximum Height ------> 1

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

but for same projectile right? and i have to determine the launch angle?

OpenStudy (anonymous):

we need the ratio of the two time of flight?

OpenStudy (anonymous):

okay, i get the question i think. thanks.

OpenStudy (maheshmeghwal9):

In first case \[\frac{u^2 \sin ^2 \theta }{2g}= \frac{u^2 \sin 2 \theta }{g}\] by solving we get \[\sin \theta = 4 \cos \theta.\] so in this case time of flight = \[\frac{2 u \sin \theta }{g}=\frac{2u \space 4 \cos \theta }{g}.\] In the second case time of flight = \[\frac{2u \sin {(90 - \theta)}}{g}= \frac{2u \cos \theta}{g.}\] So we get ratio = 4.

OpenStudy (maheshmeghwal9):

but this is nt the answer. so this is also a problem.

OpenStudy (anonymous):

the answer should be 2

OpenStudy (maheshmeghwal9):

I m also rechecking :|

OpenStudy (anonymous):

@maheshmeghwal9 hw the angle become equal in case 1

OpenStudy (anonymous):

and initila velocity

OpenStudy (anonymous):

there are diff .....i think

OpenStudy (maheshmeghwal9):

no they aren't different question is saying that in first case when body is projected with initial velocity u & angle theta then ur range = max height ok after this in second case also range & initial velocity is same but angle is different therefore angle is 90 - theta.

OpenStudy (maheshmeghwal9):

gt it?

OpenStudy (maheshmeghwal9):

gt it or nt @Yahoo! or i misunderstood the question?

OpenStudy (maheshmeghwal9):

but why i m getting 4 as answer:/

OpenStudy (anonymous):

@Ishaan94 any suggestions

OpenStudy (maheshmeghwal9):

@Ishaan94 sir plz help:) @mathslover :D

mathslover (mathslover):

not so good .. :(

OpenStudy (maheshmeghwal9):

np:) thanx for replying:)

OpenStudy (anonymous):

if range and height are same, sin^2 x = 4 sin x cos x => tan x = 4. t1 = 2u sin x/g and t2 = 2u sin y/g t2/t1 = siny/sinx, right?

OpenStudy (maheshmeghwal9):

ya but y = 90 -x

OpenStudy (anonymous):

x and y are the launch angles for first projectile and second projectile.

OpenStudy (anonymous):

y isn't 90-x? if range1 = range2 => u^2sin2x/g = u^2sin2y/g => sin2x = sin2. why isn't the answer 1?

OpenStudy (anonymous):

sin2x = sin2y*

OpenStudy (maheshmeghwal9):

no range is same thefore angles are complementary.

OpenStudy (maheshmeghwal9):

& sin x = 4 cos x

OpenStudy (maheshmeghwal9):

sin y = cos x

OpenStudy (maheshmeghwal9):

sorry i gt 4 wrong Answer is 0.25 100% sure I think options have some mistake.:/

OpenStudy (anonymous):

can anyone suggest some gud video's on the topic motion in 2-D

OpenStudy (maheshmeghwal9):

dude the solution is given by me consult ur teacher my gr8 advice answer is 0.25.

OpenStudy (anonymous):

projectile 1, speed initial = u, launch angle = x. we know x = arctan(4). projectile 2, speed initial = u, launch angle = y. if range1 = range2\[\frac{u^2\sin 2x}g =\frac{u^2\sin 2y}g \implies \sin 2x =\sin 2y \implies x=y\quad or \quad 180 - 2x = 2y\] ratio of time of flight is sinx/siny = 4 hmm seems like i am losing the grip

OpenStudy (maheshmeghwal9):

no sin y / sin x = 1/4 = 0.25 will be answer .

OpenStudy (anonymous):

.25 is same as 4.

OpenStudy (maheshmeghwal9):

ya u r right.

OpenStudy (maheshmeghwal9):

but @ yahoo options r wrong:/

OpenStudy (anonymous):

i dont think so...

OpenStudy (maheshmeghwal9):

final conclusion becoz we both gt same answers.

OpenStudy (maheshmeghwal9):

hw can u say with such believe?

OpenStudy (maheshmeghwal9):

I & @Ishaan94 both solved & gt 4 or 0.25 which one u take both are correct

OpenStudy (maheshmeghwal9):

i think the third option is 0.25 nt 2.5

OpenStudy (maheshmeghwal9):

videos on 2d @ http://www.khanacademy.org/

OpenStudy (maheshmeghwal9):

find them:)